# Learning to Evolve with Convergence Guarantee via Neural Unrolling

### v1

@@ 3 Initialization: $\bm{\omega}^{0}\leftarrow\bm{\omega}_{\text{init}}$;
     /* Step 1: Non-expansive Neural Evolution (KM Iteration) */
7     $\mathbf{d}_{IL}^{k}\leftarrow\mathbf{x}^{k-1}+\alpha\cdot\left(\mathcal{O}_{\mathrm{evo}}(\mathbf{x}^{k-1},\bm{\omega})-\mathbf{x}^{k-1}\right)$
@@ 10        $\mathbf{d}_{OL}^{k}\leftarrow\mathbf{x}^{k-1}-s_{k}\cdot\mathbf{g}_{\text{proxy}}$
    /* Step 3: Adaptive Manifold Rectification */
     $\Delta f^{k}\leftarrow f(\mathbf{d}_{OL}^{k})-f(\mathbf{d}_{IL}^{k})$;    $\bm{M}^{k}\leftarrow\sigma(-\Delta f^{k}/\tau)$
     // Soft Gating
15     $\mathbf{x}^{k}\leftarrow\text{Proj}_{\mathcal{X},\bm{P}_{\bm{\omega}}}\left(\bm{M}^{k}\odot\bm{d}_{OL}^{k}+(1-\bm{M}^{k})\odot\bm{d}_{IL}^{k}\right)$
@@
  /* Meta-Update via BPTT (Outer-Level) */
@@
18 $\nabla_{\bm{\omega}}\ell_{\mathtt{meta}}\leftarrow\textnormal{{{AutoDiff}}}_{\text{BPTT}}(\ell_{\mathtt{meta}},\{\mathbf{x}^{k}\}_{k=0}^{K})$
  $\bm{\omega}^{t+1}\leftarrow\bm{\omega}^{t}-\gamma\cdot\nabla_{\bm{\omega}}\ell_{\mathtt{meta}}$
  // Meta-Update
@@
20
21Return $\mathbf{x}^{K}$, $\bm{\omega}^{T}$
Algorithm 1 L2E: Differentiable Bilevel Evolution Strategy. The algorithm couples a neural fixed-point iteration with a dual-mode numerical operator: proxy gradients are injected as search directions during training, while inference remains purely evaluation-based.
@@ III-B Gradient-Derived Evolutionary Composite Solver
To instantiate the non-expansive operator $\mathcal{NU}(\cdot;\bm{\omega})$ introduced above, we propose a hybrid mechanism that fuses the global exploration capability of evolutionary algorithms with the local refinement power of gradient descent. This Gradient-Derived Evolutionary Composite (GEC) solver operates in a dual-mode fashion.
@@ III-B1 Learned evolutionary update
where $\mathcal{NU}$ follows the KM-style fixed-point update defined in Eq. (2), composed of the learnable operator $\mathcal{O}_{\bm{\omega}}$.
@@ III-B2 Proxy gradient update
where $s_{k}$ denotes the inner step size and $\bm{P}_{\bm{\omega}}$ is a preconditioning matrix (e.g., identity, diagonal Hessian, or low-rank approximation).
@@ III-B3 Composite update with soft-gated fusion
To seamlessly integrate these paths, we introduce an Adaptive Manifold Rectification mechanism. A dimension-wise gating mask $\mathbf{M}^{k}$ is computed based on the relative fitness improvement $\Delta f^{k}=f(\bm{d}_{OL}^{k})-f(\bm{d}_{IL}^{k})$, and apply a smooth gating function to produce a soft selection mask
@@
where $\sigma(\cdot)$ is the element-wise sigmoid function and $\tau$ is a temperature parameter controlling gating sharpness. The final update is a soft interpolation between the two candidates:
@@
where $\text{Proj}_{\mathcal{X},\bm{P}_{\bm{\omega}}}$ denotes a projection back to the feasible domain under the $\bm{P}_{\bm{\omega}}$-induced metric. This fusion mechanism balances exploitation and exploration in a differentiable, stability-preserving manner.
III-B4 Meta-Objective Function
To ensure robust meta-optimization across varying problem scales, we define $\ell_{\mathtt{meta}}$ based on the relative reduction of objective values. Let $\bar{f}(\mathbf{x})$ denote the mean objective value of a population $\mathbf{x}$. The meta-objective is formulated to minimize the negative expected normalized improvement over the task distribution $p(f)$:
@@ III-B4 Meta-Objective Function
where $\mathbf{x}^{K}(\bm{\omega})$ is the terminal population evolved on function $f$, and $\epsilon$ is a small constant for numerical stability. By minimizing this normalized loss, the framework is encouraged to maximize the proportion of cost reduction achieved within the unrolling horizon $K$, ensuring consistent optimization efficiency regardless of the absolute magnitude of the initial objective values. The complete execution protocol is formalized in Algorithm 1. This pseudocode unifies the theoretical framework into a coherent workflow: The inner loop (Lines 4-12) executes the composite update to enforce strict population contraction, while the outer loop (Lines 13-15) performs meta-optimization via BackPropagation Through Time (BPTT). By minimizing the derived meta-loss (Eq. 8), the algorithm effectively aligns the operator’s fixed point with the target optimum, bridging the gap between theoretical design and practical implementation.
@@ III-C Structured Mamba Neural Operator
To instantiate the evolution operator $\mathcal{O}_{evo}(\cdot;\omega)$, we introduce the Mamba-enhanced evolutionary operator, as illstrated in Fig. 3. Unlike sequential SSMs, we implement this as a parallelized selective block to accommodate the permutation invariance of evolutionary populations, fusing local state refinement with global exploration.
@@ III-C1 Parallel Mamba Block
The core of this operator lies in the selection mechanism, where the SSM parameters are derived directly from the input to filter noise and propagate promising signals. We transform the raw population $\mathbf{x}\in\mathbb{R}^{B\times N\times D}$ and corresponding fitness values $\mathbf{f}$ into a compact embedding $\mathbf{E}$. Then we project $\mathbf{E}$ to generate the input-dependent discretization timescale $\bm{\Delta}$ and state transformation matrices $\mathbf{B},\mathbf{C}$. We decompose the operator into two parallel streams, a state expansion stream and an input gating stream, to approximate the discretized SSM dynamics without sequential scanning. The latent state $\mathbf{s}$ and the passthrough signal $\mathbf{u}$ (i.e., $\mathbf{u}=\phi_{in}(\mathbf{E})$) are computed as:
@@
Here, $M_{\mathbf{s}}$ represents the transformed evolutionary state modulated by the selective timescale $\bm{\Delta}$. The final output of the Mamba block is regulated by a learnable multiplicative gate $\mathbf{z}$, enabling dynamic information retention:
@@
This gated formulation ($\mathbf{z}\odot M_{\mathbf{s}}+(1-\mathbf{z})\odot\mathbf{u}$) mathematically mirrors the selective retention mechanism of Gated SSMs while ensuring stable gradient flow.
@@ III-C2 Hybrid Adaptive Fusion
To complement the local refinement of Mamba with global population-level coordination, we integrate a parallel Multi-Head Self-Attention (MHSA) path: $\mathbf{H}_{attn}=\text{MHSA}(\mathbf{E})+\mathbf{E}.$ The final evolutionary proposal $\Delta_{evo}$ is synthesized via an adaptive routing mechanism that dynamically balances state exploitation and global exploration. The routing weights are determined by a lightweight non-linear router that processes population-level statistics. Specifically, a feature vector $\mathbf{s}_{pop}\in\mathbb{R}^{3}$ containing fitness standard deviation, fitness range, and population diversity is fed into an MLP to generate logits, which are then normalized:
@@
These adaptive coefficients then weigh the projected outputs of the two paths to form the final update proposal:
@@
where $\Psi_{m}$ and $\Psi_{a}$ are projection heads, and the Tanh activation ensures the proposed updates are bounded within a controlled range for stability. Crucially, to strictly satisfy the theoretical Lipschitz continuity requirement of the GNU framework, we apply Spectral Normalization to all learnable linear transformations within both the Mamba and Attention blocks. By constraining the spectral norm of weight matrices, we ensure that the global Lipschitz constant of the operator $\mathcal{O}_{evo}$ is bounded (effectively $\leq 1$), thereby guaranteeing the contractive stability of the unrolled fixed-point iteration.
@@ IV Theoretical Analysis
In this section, we establish the theoretical foundation of L2E. Specifically, we demonstrate that L2E satisfies convergence theorems under a set of regularity assumptions. In contrast to heuristic meta-optimizers, which typically lack stability guarantees, the proposed non-expansive unrolling mechanism ensures a predictable convergence trajectory. The analysis proceeds in hierarchical steps. First, we quantify the finite-time convergence rates. Second, we establish the asymptotic consistency and bilevel optimality of the framework under stronger topological conditions. Due to space constraints, we provide detailed proofs for all propositions and theorems in the Supplementary Document.
@@ IV-A Convergence Rate Analysis
    (Non-expansiveness) For any fixed $\bm{\omega}\in\Omega$, the learned composite operator $\mathcal{O}(\cdot;\bm{\omega})$ is non-expansive with respect to the population state $\mathbf{x}$: $\|\mathcal{O}(\mathbf{x};\bm{\omega})-\mathcal{O}(\mathbf{x}^{\prime};\bm{\omega})\|\leq\|\mathbf{x}-\mathbf{x}^{\prime}\|,\quad\forall\mathbf{x},\mathbf{x}^{\prime}\in\mathcal{X}.$
@@
    (Smoothness) $\ell_{\mathtt{meta}}(\mathbf{x},\bm{\omega})$ is $L_{\mathbf{x}}$-smooth w.r.t. the population $\mathbf{x}$ and $L_{\bm{\omega}}$-smooth w.r.t. the parameters $\bm{\omega}$.
@@
    (Bounded domain) The initial population $\mathbf{x}^{0}$ and $\mathbf{x}^{*}(\bm{\omega})\in\text{Fix}(\mathcal{NU}(\cdot;\bm{\omega}))$ reside within a bounded region of diameter $D_{0}$: $\|\mathbf{x}^{0}-\mathbf{x}^{*}(\bm{\omega})\|\leq D_{0}$.
@@
where $\mathbf{x}^{*}(\bm{\omega})$ is the fixed point satisfying $\mathbf{x}^{*}=\mathcal{O}(\mathbf{x}^{*};\bm{\omega})$.
@@
Under Assumptions (A1)-(A4), let the meta-parameters be updated via gradient descent: $\bm{\omega}^{t+1}=\bm{\omega}^{t}-\gamma\nabla_{\bm{\omega}}\ell_{\mathtt{meta}}(\mathbf{x}^{K}(\bm{\omega}^{t});\bm{\omega}^{t})$. With step size $\gamma<\frac{1}{L_{\bm{\omega}}+L_{\mathbf{x}}C_{K}}$, the meta-gradient norm satisfies:
@@
Remark: The bound consists of an optimization error $\mathcal{O}(1/T)$ and an approximation bias $\mathcal{O}((1-\alpha)^{2K})$, confirming that L2E consistently improves the operator towards a stationary point of the meta-loss.
IV-B Asymptotic Consistency and Bilevel Optimality
To analyze the limiting behavior of the composite solver (Algorithm 1) and the exactness of the bilevel solution, we introduce stronger topological assumptions regarding the metric space induced by the preconditioner $\mathbf{P}_{\bm{\omega}}$.
@@ IV-B Asymptotic Consistency and Bilevel Optimality
    (Closedness) The graph $\mathrm{gph}(\mathcal{O}(\cdot;\bm{\omega})):=\{(\mathbf{u},\mathbf{v})\in\mathcal{X}\times\mathcal{X}\mid\mathbf{v}=\mathcal{O}(\mathbf{u};\bm{\omega})\}$, is closed.
@@
    (Compactness and Continuity) The sets $\Omega$ and $\mathcal{X}$ are compact. For any $\bm{\omega}$, $\text{Fix}(\mathcal{NU}(\cdot;\bm{\omega}))$ is non-empty. $\ell_{\mathtt{meta}}$ is continuous on $\mathcal{X}\times\Omega$, and for fixed $\bm{\omega}$, $\ell_{\mathtt{meta}}(\cdot;\bm{\omega})$ is $L_{\ell}$-smooth, convex, and lower-bounded.
Under these structural conditions, we establish two fundamental properties: the asymptotic consistency of the inner solver and the optimality of the bilevel solution.
@@
Under Assumptions IV.2, we establish the following convergence properties for the L2E framework:
(i) Asymptotic Consistency: Under Assumptions IV.2, let $\{\mathbf{x}^{k}(\bm{\omega})\}$ be the sequence generated by Algorithm 1. For any $\bm{\omega}\in\Omega$, provided the step size satisfies $s_{k}=\frac{\kappa}{k+1}$ with $0<\kappa<\frac{\lambda_{\min}(\mathbf{P}_{\bm{\omega}})}{L_{\mathbf{x}}}$, the sequence converges to the fixed-point set, i.e., $\lim_{k\to\infty}\mathrm{dist}(\mathbf{x}^{k}(\bm{\omega}),\mathrm{Fix}(\mathcal{NU}(\cdot;\bm{\omega})))=0$, and the meta-objective value converges to the optimum:
@@
(ii) Bilevel Optimality: As the unrolling depth $K\to\infty$, let $\{(\mathbf{x}^{K}(\bm{\omega}^{K}),\bm{\omega}^{K})\}$ be the sequence of generated solutions. We prove that any limit point $(\bar{\mathbf{x}},\bar{\bm{\omega}})$ constitutes a valid solution to the fixed-point constrained bilevel problem (Eq. 1), satisfying $\bar{\bm{\omega}}\in\arg\min_{\bm{\omega}\in\Omega}\varphi(\bm{\omega})$ and $\bar{\mathbf{x}}=\mathcal{NU}(\bar{\mathbf{x}};\bar{\bm{\omega}})$. Furthermore, the sequence of surrogate objectives converges to the true optimum, i.e., $\lim_{K\to\infty}\inf_{\bm{\omega}\in\Omega}\varphi_{K}(\bm{\omega})=\inf_{\bm{\omega}\in\Omega}\varphi(\bm{\omega})$.

### v2

@@ 3 Initialization: $\bm{\omega}^{0}\leftarrow\bm{\omega}_{\text{init}}$;
     /* Step 1: Stability-inducing Neural Evolution (KM Iteration) */
7     $\mathbf{d}_{IL}^{k}\leftarrow(1-\alpha)\mathbf{x}^{k-1}+\alpha\cdot\mathcal{O}_{\mathrm{evo}}(\mathbf{x}^{k-1},\bm{\omega})$
@@ 10        $\mathbf{d}_{OL}^{k}\leftarrow\mathbf{x}^{k-1}-s_{k}\cdot\mathbf{g}_{\text{proxy}}$
    /* Step 3: Soft-gated Fusion */
15     $\Delta f^{k}\leftarrow f(\mathbf{d}_{OL}^{k})-f(\mathbf{d}_{IL}^{k})$;    $\bm{M}^{k}\leftarrow\sigma(-\Delta f^{k}/\tau)$ $\mathbf{x}^{k}\leftarrow\text{Proj}_{\mathcal{X},\bm{P}_{\bm{\omega}}}\left(\bm{M}^{k}\odot\bm{d}_{OL}^{k}+(1-\bm{M}^{k})\odot\bm{d}_{IL}^{k}\right)$
@@
  /* Meta-Update via AutoDiff (Outer-Level) */
@@
 $\nabla_{\bm{\omega}}\ell_{\mathtt{meta}}\leftarrow\textnormal{{{AutoDiff}}}(\ell_{\mathtt{meta}},\{\mathbf{x}^{k}\}_{k=0}^{K})$
  // BPTT
18  $\bm{\omega}^{t+1}\leftarrow\bm{\omega}^{t}-\gamma\cdot\nabla_{\bm{\omega}}\ell_{\mathtt{meta}}$
@@
20Return $\mathbf{x}^{K}$, $\bm{\omega}^{T}$
Algorithm 1 L2E: Differentiable Bilevel Evolution Strategy. The algorithm couples a KM-style averaged neural update with dual-mode guidance: proxy gradients are injected during training, while inference remains evaluation-only.
@@ III-B Gradient-derived Evolutionary Composite Solver
To instantiate the proposed operator $\mathcal{NU}(\cdot;\bm{\omega})$ introduced above, we propose a hybrid mechanism that fuses the global exploration capability of evolutionary algorithms with the local refinement power of gradient descent. This GEC solver operates in a dual-mode fashion.
@@ III-B1 Learned evolutionary update
where $\mathcal{NU}$ follows the KM-style fixed-point update defined in Eq. (2), which is composed of the learnable operator $\mathcal{O}_{\bm{\omega}}$.
@@ III-B2 Proxy Gradient Update
where $s_{k}$ denotes the inner step size, and $\bm{P}_{\bm{\omega}}$ represents a preconditioning matrix, such as an identity matrix, a diagonal Hessian, or a low-rank approximation.
@@ III-B3 Composite update with soft-gated fusion
A dimension-wise gating mask $\mathbf{M}^{k}$ is computed based on the relative fitness improvement $\Delta f^{k}=f(\bm{d}_{OL}^{k})-f(\bm{d}_{IL}^{k})$, and apply a smooth gating function to produce a soft selection mask
@@
where $\sigma(\cdot)$ is the element-wise sigmoid function and $\tau$ is a temperature parameter that controls the gating sharpness. The final update is computed as a soft interpolation between the two candidates:
@@
where $\text{Proj}_{\mathcal{X},\bm{P}_{\bm{\omega}}}$ denotes a projection back onto the feasible set under the $\bm{P}_{\bm{\omega}}$-induced metric. [^1] The subscript $(\mathcal{X},\mathbf{P}_{\bm{\omega}})$ indicates a metric projection onto the feasible set $\mathcal{X}$ under the $\mathbf{P}_{\bm{\omega}}$-induced norm: $\mathrm{Proj}_{\mathcal{X},\mathbf{P}_{\bm{\omega}}}(\mathbf{z}):=\arg\min_{\mathbf{x}\in\mathcal{X}}\ \|\mathbf{x}-\mathbf{z}\|_{\mathbf{P}_{\bm{\omega}}}^{2},\qquad\|\mathbf{v}\|_{\mathbf{P}_{\bm{\omega}}}^{2}:=\mathbf{v}^{\top}\mathbf{P}_{\bm{\omega}}\mathbf{v}.$ This fusion mechanism balances exploitation and exploration. It operates in a manner that is both differentiable and stability-preserving.
III-B4 Meta-objective
Let $\bar{f}(\mathbf{x})$ denote the mean objective value of a population $\mathbf{x}$. We define the meta-objective as the negative expected normalized improvement over the task distribution $p(f)$:
@@ III-B4 Meta-objective
where $\mathbf{x}^{K}(\bm{\omega})$ is the terminal population after $K$ unrolled updates on function $f$, and $\epsilon$ is a small constant for numerical stability. This normalization makes the meta-loss less sensitive to the absolute scale of objective values, enabling consistent training across tasks with different magnitudes. In Algorithm 1, we minimize $\ell_{\mathtt{meta}}$ via BPTT to update $\bm{\omega}$, thereby shaping the learned evolutionary dynamics to improve terminal performance under a fixed evaluation budget.
@@ III-C Structured Mamba Neural Operator
To instantiate the evolution operator $\mathcal{O}_{\mathrm{evo}}(\cdot;\bm{\omega})$, we adopt a structured Mamba-based operator implemented in a parallel selective form to respect the permutation-invariant nature of populations (Fig. 3).
@@ III-C1 Parallel Mamba Block
The operator takes the population $\mathbf{x}\in\mathbb{R}^{B\times N\times D}$ and fitness values $\mathbf{f}$ as input and maps them to an embedding $\mathbf{E}$. We then generate input-dependent parameters, including the discretization timescale $\bm{\Delta}$ and state transformation matrices $\mathbf{B},\mathbf{C}$, and compute two parallel streams: a state-expanded representation and a passthrough signal. Specifically,
@@
The output is computed via a gated combination:
@@ III-C2 Hybrid Adaptive Fusion
To incorporate population-level coordination, we add a parallel multi-head self-attention path $\mathbf{H}_{attn}=\text{MHSA}(\mathbf{E})+\mathbf{E}.$ The final evolutionary proposal is obtained by routing between the two paths using population statistics $\mathbf{s}_{pop}$:
@@
Here, $\Psi_{m}$ and $\Psi_{a}$ denote projection heads. The Tanh activation bounds the update magnitude, serving as a stability-oriented design choice. Here we apply Spectral Normalization [48] to the learnable linear transformations in both paths to control update sensitivity. Together, spectral normalization and bounded updates act as stability-oriented regularization, making the learned updates less sensitive to input perturbations in practice.
@@ IV Theoretical Analysis
This section provides a stability-oriented theoretical analysis for L2E. We do not claim universal guarantees for arbitrary neural components. Instead, we clarify the sufficient conditions under which the unrolled operator iteration exhibits controlled dynamics and admits meaningful optimization bounds. Due to space constraints, detailed proofs are deferred to the Supplementary Document.
@@ IV-A Convergence Rate Analysis
    (Operator regularity) For any fixed $\bm{\omega}\in\Omega$, the learned composite operator $\mathcal{O}(\cdot;\bm{\omega})$ is non-expansive with respect to the population state $\mathbf{x}$: $\|\mathcal{O}(\mathbf{x};\bm{\omega})-\mathcal{O}(\mathbf{x}^{\prime};\bm{\omega})\|\leq\|\mathbf{x}-\mathbf{x}^{\prime}\|,\ \forall\mathbf{x},\mathbf{x}^{\prime}\in\mathcal{X}.$
@@
    (Smoothness) $\ell_{\mathtt{meta}}(\mathbf{x},\bm{\omega})$ is $L_{\mathbf{x}}$-smooth with respect to the population $\mathbf{x}$ and $L_{\bm{\omega}}$-smooth with respect to the parameters $\bm{\omega}$.
@@
    (Bounded domain) The initial population $\mathbf{x}^{0}$ and $\mathbf{x}^{*}(\bm{\omega})\in\text{Fix}(\mathcal{NU}(\cdot;\bm{\omega}))$ are contained within a bounded region of diameter $D_{0}$: $\|\mathbf{x}^{0}-\mathbf{x}^{*}(\bm{\omega})\|\leq D_{0}$.
@@
where $\mathbf{x}^{*}(\bm{\omega})$ is the fixed point that satisfies $\mathbf{x}^{*}=\mathcal{O}(\mathbf{x}^{*};\bm{\omega})$.
@@
Under Assumptions (A1)-(A4), let the meta-parameters be updated via gradient descent: $\bm{\omega}^{t+1}=\bm{\omega}^{t}-\gamma\nabla_{\bm{\omega}}\ell_{\mathtt{meta}}(\mathbf{x}^{K}(\bm{\omega}^{t});\bm{\omega}^{t})$. Provided that the step size is $\gamma<\frac{1}{L_{\bm{\omega}}+L_{\mathbf{x}}C_{K}}$, the meta-gradient norm satisfies:
@@
Remark: The bound decomposes into an optimization term $\mathcal{O}(1/T)$ and an approximation term $\mathcal{O}((1-\alpha)^{2K})$. This decomposition explicitly characterizes the influence of the unrolling depth $K$ and the averaging factor $\alpha$ on the meta-update quality under Assumptions (A1)-(A4). Moreover, the analysis aligns with established bilevel optimization literature, in which guarantees for convergence and consistency are derived from operator regularity and topological conditions [52, 36].
IV-B Asymptotic Consistency under Fixed-point Constraints
To interpret the limiting behavior of Algorithm 1 under the fixed-point constrained formulation, we consider stronger (idealized) topological assumptions. These assumptions facilitate an asymptotic consistency analysis. Specifically, this analysis connects the unrolled iteration to the fixed-point set of Eq. (1) and elucidates its relationship to the induced outer objective.
@@ IV-B Asymptotic Consistency under Fixed-point Constraints
    (Closedness) The graph $\mathrm{gph}(\mathcal{O}(\cdot;\bm{\omega})):=\{(\mathbf{u},\mathbf{v})\in\mathcal{X}\times\mathcal{X}\mid\mathbf{v}=\mathcal{O}(\mathbf{u};\bm{\omega})\}$ is closed.
@@
    (Compactness and Continuity) The sets $\Omega$ and $\mathcal{X}$ are compact. Furthermore, for any $\bm{\omega}$, the set $\text{Fix}(\mathcal{NU}(\cdot;\bm{\omega}))$ is non-empty. Additionally, $\ell_{\mathtt{meta}}$ is continuous on $\mathcal{X}\times\Omega$. Finally, for a fixed $\bm{\omega}$, the function $\ell_{\mathtt{meta}}(\cdot;\bm{\omega})$ is $L_{\ell}$-smooth, convex, and lower-bounded.
@@
Based on Assumption IV.2, we establish the following convergence properties for L2E:
(i) Let $\{\mathbf{x}^{k}(\bm{\omega})\}$ denote the sequence generated by Algorithm 1. For any $\bm{\omega}\in\Omega$, assume the step size satisfies $s_{k}=\frac{\kappa}{k+1}$, where $0<\kappa<\frac{\lambda_{\min}(\mathbf{P}_{\bm{\omega}})}{L_{\mathbf{x}}}$. Under these conditions, the sequence converges to the fixed-point set, i.e., $\lim_{k\to\infty}\mathrm{dist}(\mathbf{x}^{k}(\bm{\omega}),\mathrm{Fix}(\mathcal{NU}(\cdot;\bm{\omega})))=0$. Consequently, the meta-objective value converges to the optimum:
@@
(ii) Let $\{(\mathbf{x}^{K}(\bm{\omega}^{K}),\bm{\omega}^{K})\}$ denote the sequence of generated solutions as the unrolling depth $K\to\infty$. We define the induced outer objective (value function) as $\varphi(\bm{\omega}):=\inf_{\mathbf{x}\in\mathrm{Fix}(\mathcal{NU}(\cdot;\bm{\omega}))}\ \ell_{\mathtt{meta}}(\mathbf{x};\bm{\omega}),$ $\varphi_{K}(\bm{\omega}):=\ell_{\mathtt{meta}}(\mathbf{x}^{K}(\bm{\omega});\bm{\omega}).$ We prove that any limit point $(\bar{\mathbf{x}},\bar{\bm{\omega}})$ is a valid solution to the fixed-point constrained bilevel problem (Eq. (1)). In particular, the limit point satisfies $\bar{\bm{\omega}}\in\arg\min_{\bm{\omega}\in\Omega}\varphi(\bm{\omega})$ and $\bar{\mathbf{x}}=\mathcal{NU}(\bar{\mathbf{x}};\bar{\bm{\omega}})$. Furthermore, the sequence of surrogate objectives converges to the optimum of the induced outer objective, i.e., $\lim_{K\to\infty}\inf_{\bm{\omega}\in\Omega}\varphi_{K}(\bm{\omega})=\inf_{\bm{\omega}\in\Omega}\varphi(\bm{\omega})$.